# Daco: Towards Application-Driven and Comprehensive Data Analysis via Code Generation

### v1

@@ Abstract
Data analysis is a crucial analytical process to generate in-depth studies and conclusive insights to comprehensively answer a given user query for tabular data. In this work, we aim to propose new resources and benchmarks to inspire future research on this crucial yet challenging and under-explored task. However, collecting data analysis annotations curated by experts can be prohibitively expensive. We propose to automatically generate high-quality answer annotations leveraging the code-generation capabilities of LLMs with a multi-turn prompting technique. We construct the Daco dataset, containing (1) 440 databases (of tabular data) collected from real-world scenarios, (2) query-answer pairs that can serve as weak supervision for model training, and (3) a concentrated but high-quality test set with human refined annotations that serves as our main evaluation benchmark. We train a 6B supervised fine-tuning (SFT) model on Daco dataset, and find that the SFT model learns reasonable data analysis capabilities. To further align the models with human preference, we use reinforcement learning to encourage generating analysis perceived by human as helpful, and design a set of dense rewards to propagate the sparse human preference reward to intermediate code generation steps. Our Daco-RL algorithm is evaluated by human annotators to produce more helpful answers than SFT model in cases, validating the effectiveness of our proposed algorithm. Data and code are released at https://github.com/shirley-wu/daco.
@@ 1 Introduction
Data analysis is the process of systematically applying statistical and/or logical reasoning to evaluate and comprehend data. Existing literature has investigated answering queries about information given by structural data (e.g., tables) (Chen et al., 2021a; Nan et al., 2022; Lu et al., 2023). However, they either focus on straightforward factual retrieval or short-form entity/arithmetic resolutions for specifically given entities, while real-world data analysis can involve more complex analytical processes.
@@
Take the scenario in Figure 1 as an example: a user is investigating potential age discrimination of a shop. To effectively answer queries such as this one, a chain of mathematical and logical reasoning and interacting with the data is required. For instance, finding 1 is inferred from analyzing age distribution within the membership data (‘member’ table), while finding and suggestion 2 are derived by comparing the participants’ ages during the happy hours (using both ‘member’ and ‘happy_hour_member’ tables). These rigorous quantitative analyses eventually conclude the opposite to the user’s hypothesis. As valuable as the conclusive suggestions such comprehensive analysis can bring, the extensive labor-efforts, hinted by these examples, can hinder the efficiency of gaining intelligence from the data in a competitive business environment. It is thus imperative to devise a system that is able to automate the aforementioned data analysis process.
@@
To this end, we introduce a new dataset for this challenging task, Daco, data analysis via code generation. Daco is constructed from a set of diverse real-world databases associated with curated user queries. In light of the previously described labor-intensive challenge, we propose to leverage LLMs with a multi-turn chained prompts to automatically curate the analytical answers for each query. Specifically, our designed framework employs the code generation capabilities of GPT-4 (OpenAI, 2023) for automating the statistical analysis, interleaved with its ability to interpret the obtained quantitative results. The Daco dataset contains databases and associated user queries, which can be used for both model fine-tuning and evaluation. To provide a refined benchmarking resource, we curate a high-quality test set through comprehensive human annotations on a subset of 100 samples. Detailed statistics are in Table 1.
Although LLM exhibit reasonable analytical capabilities (and hence we are able to automatically curate the pre-refined data), we empirically find that its generations often fall short of human expectations of what good analyses should be (e.g., relevance to the queries, logical coherency, and higher-level quantitative interpretations). To further improve the generations by aligning the models with corresponding human preferences, we design a reinforcement learning algorithm (Daco-RL) that leverages two newly designed reward models (RM), i.e., contribution RM and regularization RM, to efficiently provide denser feedback. Concretely, the contribution RM heuristically provides better learning signals for the intermediate code generation steps (for more relevant quantitative analysis), while regularization RM helps preventing reward hacking (Skalse et al., 2022) of typical RLHF models (Casper et al., 2023). We test our algorithm on a fine-tuned CodeGeeX-6B model (Zheng et al., 2023a), where the win rate of in human-annotated pairwise comparison justifies its effectiveness on learning to generate human preferred analyses.
@@
In summary, our contributions are three folds: (1) We explore the challenging task of data analysis, where we construct the Daco dataset with our proposed multi-turn prompting technique on a diverse set of real-world databases. (2) We curate a human-refined evaluation set for benchmarking models. (3) We design the Daco-RL algorithm to jointly optimize code generation and answer generation towards human alignment, which demonstrates a significant human evaluated win rate on the helpfulness metric.
2 The Daco Task and Dataset
@@
As shown in Figure 1, the input to our task is a database and a query , and the output answer is formatted as two lists of findings and suggestions respectively. In this work, the database should be a relational database containing multiple named tables.
We construct our Daco dataset through four stages: (1) database collection, (2) query collection, (3) automatic annotation collection, and (4) human refinement. The workflow is illustrated in Figure 2. Our final dataset contains the training, development and test sets with annotations generated by GPT-4, along with a human-refined testing subset. To distinguish the two test sets, we use Test to represent the automatically annotated set and Test to represent the human refined one. Statistics are shown in Table 1.
Database collection. We collect databases from two sources: Spider (Yu et al., 2018) and Kaggle (https://www.kaggle.com/datasets). There are 157 databases collected from Spider, which originally come from university databases, DatabaseAnswers and Wikipedia. We additionally crawl and filter 5,830 databases from Kaggle. From this pool, we manually select a subset of 314 clean and interpretable databases to build our dataset. To maintain the diversity of the resulting database set, 157 of the databases are deliberately chosen near the long tail of its topic distribution. For this, we employ BERTopic (Grootendorst, 2022) to model the topic distribution, which produces in total 160 topics. We take its least frequent 80 topics as the long tail, which covers 26.79% of the total databases.
In total, Daco comprises 471 databases, each of which contains on average 2.3 tables. To better visualize the major topic distribution of this selected subset, we again use BERTopic but group these databases into 10 topics. The keywords for top 5 topics are shown in Figure 4. The leading topic (topic 1) is associated with business setting and consists of 46.52% of the dataset. The remaining nine topics exhibit a relatively even distribution, covering a broad range of domains, including sports (topic 2), healthcare (topic 3), weather (topic 4), and education (topic 5).
Query collection. We generate 10 queries for each database by prompting ChatGPT to first assume the role of a database stakeholder and then generate an application-driven query based on the role. To ensure the quality of the query, we perform a manual filtering to the machine generated queries. Specifically, we remove queries that are not driven by real-world applications or cannot be answered by the given reference database. We train a group of 6 annotators to perform such a filtering process. As a result, there are about 42% of the queries removed, where the removal agreement achieves a 0.62 cohen kappa score.
After the aforementioned processes, we obtain in total 2,664 queries. We show the top 15 verbs and their top 3 direct noun objectives in Figure 4. The queries demonstrate a notable level of diversity. The most common type of queries is to request analysis (such as “analyze data” and “identify pattern”), followed by queries aiming to make decisions (such as “determine strategy” and “make decision”).
Automatic annotation collection. As shown in the right half of Figure 5, we design a pipeline that leverages the code generation capability of LLMs to automate the answer annotation for our Daco dataset. Based on the database and the query, we instruct the LLM to perform data analysis in multiple turns. At each turn, the LLM will produce a python code snippet and take its execution outputs as evidences to reason over and support its follow-up interpretation. After each turn, we prompt the model to decide whether the analysis is sufficiently comprehensive; if deemed sufficient, it terminates the coding turns and produces the final answer.
With this pipeline, we instruct GPT-4 to automatically generate all the answer annotations to each query of our dataset, for both the intermediate code and the final analysis answering the queries. To improve the quality of such automatically constructed annotations, we additionally allow GPT-4 to correct its own mistakes when its generated code leads to run-time or syntax error, where only the corrected codes are kept. In total, we obtain 1.9k valid query-answer pairs, each with roughly 3.3 intermediate coding steps.
Human refinement. The annotated analyses thus far have been algorithmically generated, where their actual quality are to be further verified. We thus curate a human-refined subset containing 100 densely human-annotated query-answer pairs. For each query, we sample 3 different analysis candidates using the previously described automated method (with GPT-4). We ask the annotators to evaluate the quality of each machine generated bullet point and categorize each point into one of bad, borderline, or good (associated with scalar scores of 0, 1 and 2 respectively, the higher the better) to each. The bullet points deemed higher quality are then mixed (from the 3 candidate analyses) and refined (with a few manual textual edits) into one final gold-analysis. In the refinement stage, the annotators should first combine all bullet points ranked as “good”, remove duplicate points, and reorder the points to maintain a coherent flow. Suppose the number of bullet points are lower than our pre-defined lowest threshold (3 bullet points per answer), the annotators should select bullet points ranked as “borderline” to augment the answer. we ask a group of 3 internal members to perform refinement. The agreement accuracy of the refinement process (candidate point selection) is 0.83 and the Cohen’s Kappa is 0.67.
@@
Evaluation. To evaluate the quality of generated data analysis, we use helpfulness as the main metric. Motivated by literature in the data analysis field (Long & Long, 2009), we define helpfulness as: (1) relevance to the query, (2) effective and insightful data interpretation, and (3) diversity in terms of analysis perspectives. We evaluate helpfulness through pairwise comparison following common approach (Ouyang et al., 2022; Wu et al., 2023; Zheng et al., 2023b). Given two analyses generated by two different systems, the annotator (either human or simulated by ChatGPT) selects the more helpful one based on our defined criteria. The winning rate of each system is reported as helpfulness score. To obtain a comparable set of numbers for all models, we report the winning rate of each model against Test and Test annotations. The upper bound for this score would be 50, as a score of 50 indicates that the model generations are perceived as helpful as annotations.
Code generation (right): At each turn, the model generates python code, execute the code with a code interpreter, and reads the execution outputs. Eventually, it summarizes the results into a final answer.
Daco-RL (left): The answer RM evaluates the helpfulness of the final answer and is our end optimization goal. However, to provide denser reward, we use a contribution RM to evaluate how much each generated code contribute to the final answer. The reward is provided at the end of each generated code snippet. Contribution RM is vulnerable to reward hacking, which motivates us to propose regularization RM to discourage reward hacking generations.
3 Daco-RL
While Daco contains mostly algorithmic machine generated analyses, the machine generations without human refinement cannot well align with human preferences (of “good” analyses). Our human refinement process shows that only 47.4% bullet points are “good” points perfectly addressing user queries; the majority of 52.2% are evaluated as “borderline” points that only partially aligns with human expectations; and the remaining 0.4% are considered as “bad”.
We are therefore interested in investigating whether aligning human preferences via an RLHF fashion could lead to better machine generated analyses. We thus propose the Daco-RL algorithm, which is illustrated in the left half of Figure 5. Our end goal is to optimize the helpfulness of the analyzed points, which is modelled with an answer RM . In addition to this sparse reward signal, we use a heuristically defined contribution RM to reward each intermediate step, which is further regularized with a regularization RM to prevent reward hacking. In the following sections, we first explain the three reward models sequentially, and eventually explain our whole RLHF pipeline.
Notations. We train a language model that interacts with the python interpreter in a conversational manner. Formally, the full dialogue is a list of messages , where , , and stand for human message, code, execution outputs and final answer respectively. The dialogue starts with human message containing both query and database meta-data. In the later messages, code and final answer are generated by our model, while execution outputs are produced by the python interpreter. To feed the dialogue into the language model, we wrap each message between a begin-of-message phrase (denoted as BOM) and the end-of-sentence token, and then concatenate all messages together. We use a different BOM for each type of message. Thus, at each turn of the conversation, the language model can decide whether to generate code or final answer by generating the corresponding BOM.
Answer RM. Our end goal is to optimize the helpfulness of answer , which is modelled by the answer RM . Particularly, we model the helpfulness of each single bullet point rather than the full answer. We collect pairwise comparison data of bullet points from ChatGPT to train . Given a pair of bullet points where one is preferred over the other, is trained to assign a higher score to the preferred bullet point. Given access to the full conversation, produces a reward score at the end of each bullet point in the answer.
To encourage diversity, we additionally add a repetition penalty. Given a list of findings , we encourage the -th finding to be different from all previous findings by imposing a penalty score as , where computes the similarity between two bullet points.111 is computed as the cosine similarity of Sentence-BERT (Reimers & Gurevych, 2019) embeddings. We use the all-MiniLM-L6-v2 model. This score is then subtracted from the reward score from . The same procedure is applied to the suggestion list.
Contribution RM. The answer RM described above is a sparse reward signal that rewards only the answer but not intermediate coding steps . To provide denser reward signal for optimizing , we aim to evaluate the helpfulness of each coding step . However, annotating helpfulness of intermediate steps is much more difficult. For one, the helpfulness of python code is more vague to define, and the other, evaluating helpfulness requires coding expertise, which makes it more expensive to collect human annotations.
To measure the helpfulness of intermediate steps without the huge expense required by human annotations, we heuristically define the helpfulness as how much an intermediate step contributes to the final answer. Concretely, we compute the similarity between final answer and code outputs to measure the helpfulness of . We use to rank the helpfulness of different steps, and use the comparison pairs between intermediate steps to train the contribution RM . Given the conversation, predicts the contribution level of each as its reward score. does not take the execution output into consideration when scoring , which simplifies model implementation and excludes spurious correlation the model may exploit from the surface form of .
Regularization RM. The heuristically defined contribution RM may not necessarily perfectly align with the true helpfulness. This is due to a known reward misspecification issue termed reward hacking (Skalse et al., 2022), where the policy model achieves higher scores from the reward model but its true reward decreases. We propose to regularize such behavior with a regularization RM . Given the misspecified reward model, in our case, we first train an RL model until its generations start to collapse to certain patterns. These patterns typically receive high rewards from but do not align well with human expectation, and thus are considered as reward hacking behaviors. We denoted this RL model as . We use to produce generations with typical reward hacking behaviors. These generations are paired with generations without reward hacking behaviors, such as generations from supervised fine-tuning (SFT) model or the pre-human refined answers generated from GPT-4, to further train the regularization RM . As an intuition, this means will assign lower scores to typical reward hacking behaviors.
RLHF. For our whole Daco-RL pipeline, we optimize the language model against the mixture of all three aforementioned rewards, , and .
More specifically, we first train a multi-task reward model denoted as to jointly learn and . This guarantees that the reward score distribution for and are relatively close, so the reward signal for and will not overshadow each other. We train a separate model to learn , and mix and into when rewarding .222In practice, we find has a large non-zero mean value, so we subtract it before weighted average. Here, is a hyper-parameter to balance the variance of and tuned to maximize reward model accuracy on development set. This guarantees the mixed reward can both encourage high contribution and penalize reward hacking.
We use proximal policy optimization (PPO) (Schulman et al., 2017) as our learning algorithm. During training, PPO jointly optimizes a value model and a policy model . The objective of the policy model is to optimize the generalized advantage estimation (Schulman et al., 2016) , where for each time step . When applied to text generation, the generative language model is the policy model and each generated token is an action. In our multi-turn conversational setting, however, only part of the tokens in the dialogue are generated by language model (concretely, and ). In other words, although the language model still takes the full conversation as input, we only compute GAE and gradients over the model generated subsequence, i.e., .

### v2

@@ Abstract
Data analysis is a crucial analytical process essential for deriving insights from real-world databases. As shown in Figure 1, the need for data analysis typically arises from specific application scenarios, and requires diverse reasoning skills including mathematical reasoning, logical reasoning, and strategic reasoning. Existing work often focus on simple factual retrieval or arithmetic resolutions and thus are insufficient for addressing complex real-world queries. This work aims to propose new resources and benchmarks on this crucial yet challenging and under-explored task. Due to the prohibitively high cost of collecting expert annotations, we use large language models (LLMs) enhanced by code generation to automatically generate high-quality data analysis, which will later be refined by human annotators. We construct the Daco dataset, containing (1) 440 databases (of tabular data) collected from real-world scenarios, (2) automatically generated query-answer pairs that can serve as weak supervision for model training, and (3) a concentrated but high-quality test set with human refined annotations that serves as our main evaluation benchmark. Experiments show that while LLMs like GPT-4 exhibit promising data analysis capabilities, they are still evaluated as less helpful than human-written analysis on 58.1% cases. Leveraging our weak supervision data, we experiment with various fine-tuning methods, including supervised fine-tuning (SFT) and reinforcement learning from human feedback (RLHF). Our trained model outperforms existing baselines for table question answering, and RLHF further boosts the helpfulness of generated analysis on 58.5% cases. Data and code are released at https://github.com/shirley-wu/daco.
@@ 1 Introduction
Data analysis is the process of systematically applying statistical and logical reasoning to comprehend data and derive insights. Existing literature on table question answering has investigated answering queries about information given by structural data (e.g., tables) [7, 26, 24]. However, they either focus straightforward factual retrieval or short-form arithmetic resolutions over retrieved entities while real-world data analysis can involve more complex analytical processes.
@@
Take the scenario in Figure 1 as an example: a user is investigating potential age discrimination of a shop. To effectively answer queries such as this one, a chain of mathematical and logical reasoning and interacting with the data is required. For instance, finding 1 is inferred from the membership data (‘member’ table) through mathematical and analytical reasoning, while finding and suggestion 2 are derived from both ‘member’ and ‘happy_hour_member’ tables through mathematical and strategic reasoning. These rigorous quantitative analyses eventually conclude the opposite to the user’s hypothesis. As valuable as the conclusive suggestions such comprehensive analysis can bring, the extensive labor-efforts, hinted by these examples, can hinder the efficiency of gaining intelligence from the data in a competitive business environment. It is thus imperative to devise a system that is able to automate the aforementioned data analysis process.
@@
To this end, we introduce a new dataset for this challenging task, Daco, data analysis via code generation. Daco is constructed from a set of diverse real-world databases associated with curated user queries. In light of the previously described labor-intensive challenge, we propose to leverage LLMs with a multi-turn chained prompts to automatically curate the analytical answers for each query. Specifically, our designed framework employs the code generation capabilities of GPT-4 [28] for automating the statistical analysis, interleaved with its ability to interpret the obtained quantitative results. The Daco dataset contains databases and associated user queries, where each query is annotated with an average of coding steps and lines of code during intermediate steps and a final output of bullet points. This resource can be used for both model fine-tuning and evaluation. To provide a refined benchmarking resource, we curate a high-quality test set through comprehensive human annotations on a subset of 100 samples. Detailed statistics are in Table 1.
We evaluate three types of methods on this new dataset: (1) existing methods designed for table QA tasks, (2) prompt-based LLMs, and (3) fine-tuned LLMs. We use helpfulness as our main evaluation metric, assessed through pair-wise comparison. We observe that proprietary LLMs such as GPT-4 demonstrate strong data analysis capabilities and significantly outperform table QA models. However, they still fall short of human performance by 58.1% in pair-wise evaluations. Regarding fine-tuned LLMs, via supervised fine-tuning (SFT) using automatically generated annotations that include both code generation trajectories and the final answers. Inspired by the recent success of reinforcement learning from human feedback (RLHF) [29, 35, 4, 5, 48, 38], we employ RLHF to further align the SFT models with human preferences towards helpful data analysis. Our SFT model exhibits promising data analyais capabilities and outperforms table QA baselines despite lagging behind proprietary LLMs. On top of SFT, RLHF further enhances the output helpfulness in 58.5% of cases.
@@
In summary, our contributions are as follows: (1) We explore the challenging task of data analysis, where we construct the Daco dataset with our proposed multi-turn prompting technique on a diverse set of real-world databases. (2) We curate a human-refined evaluation set for benchmarking models. (3) We evaluate a diverse set of models on this challenging dataset, including existing models designed for table QA, prompt-based LLMs, and fine-tuned LLMs with both SFT and RLHF. (4) Our dataset and code are made publicly available at https://github.com/shirley-wu/daco.
2 Task Formulation
@@
As shown in Figure 1, the input to our task consists of a database and a query , where the database is a relational database containing multiple named tables. The output is formatted as two lists: findings and suggestions.
Inspired by recent work on tool usage and LLM agent [40, 41, 23], we allow the LLM to invoke tools for multiple turns before producing the final output . At the -th turn, the action is generated by the LLM, and the observation is produced by the environment after executing the corresponding action . While the action can involve various tools such as SQL executor or search engine, we utilize Python due to its extensive mathematical libraries and programming flexibility.
@@
To evaluate the quality of generated data analysis , we use helpfulness as the primary metric. Motivated by literature in the data analysis field [22], we define helpfulness as: (1) relevance to the query, (2) effective and insightful data interpretation, and (3) diversity in terms of analysis perspectives. We evaluate helpfulness through pairwise comparison following common approach [29, 38, 45]. Given two analyses generated by two different systems, the annotator (either human or simulated by LLM) selects the more helpful one based on our defined criteria. The winning rate of each system is reported as helpfulness score. To obtain a comparable set of numbers for all models, we report the winning rate of each model against the annotations, so a score of 50 would indicate the model generations are perceived as helpful as annotations.
3 Dataset Construction
We construct our Daco dataset through four stages: (1) database collection, (2) query collection, (3) automatic annotation collection, and (4) human refinement. The workflow is shown in Figure 2(a).
Database collection. We collect databases from two sources: Spider [42] and Kaggle (https://www.kaggle.com/datasets). There are 157 databases collected from Spider, which originally come from university databases, DatabaseAnswers and Wikipedia. We then crawl and filter 5,830 databases from Kaggle. Despite the initial filtering, the quality of these databases remains low, often containing noisy or unintelligible data. Therefore, we conduct a manual inspection and thereby select a subset of 314 clean and interpretable databases to build our dataset. To maintain the diversity of the resulting database set, 157 of the databases are deliberately chosen from the long tail of the topic distribution. We employ BERTopic [11] to model the topic distribution, which produces in total 160 topics. We take its least frequent 80 topics as the long tail, which covers 26.79% of the total databases.
Query collection. We generate 10 queries for each database by prompting ChatGPT to first assume the role of a database stakeholder and then generate an application-driven query based on the role. To ensure the quality of the query, we perform a manual filtering to the machine generated queries. Specifically, we remove queries that are not driven by real-world applications or cannot be answered by the given reference database. We train a group of 6 annotators to perform such a filtering process. As a result, there are about 42% of the queries removed, where the removal agreement achieves a 0.62 cohen kappa score. After the aforementioned processes, we obtain in total 2,664 queries. Examples of databases and the automatically generated queries are shown in Figure 8.
Automatic annotation collection. As shown in Figure 2(b), we design a pipeline that leverages the code generation capability of LLMs to automate the answer annotation for our Daco dataset. Based on the database and the query, we instruct the LLM to perform data analysis in multiple turns. At each turn, the LLM will produce a python code snippet and take its execution outputs as evidences to reason over and support its follow-up interpretation. After each turn, we prompt the model to decide whether the analysis is sufficiently comprehensive; if deemed sufficient, it terminates the coding turns and produces the final answer. With this pipeline, we instruct GPT-4 to automatically generate all the answer annotations to each query of our dataset, for both the intermediate code and the final analysis answering the queries. To improve the quality of such automatically constructed annotations, we additionally allow GPT-4 to correct its own mistakes when its generated code leads to run-time or syntax error, where only the corrected codes are kept. In total, we obtain 1.9k valid query-answer pairs, each with roughly 3.3 intermediate coding steps.
Human refinement. The annotated analyses thus far have been algorithmically generated, where their actual quality are to be further verified. We thus curate a human-refined subset containing 100 densely human-annotated query-answer pairs as illustrated in Figure 2(c). For each query, we sample
3 different analysis candidates using the previously described automated method with GPT-4. We ask the annotators to evaluate the quality of each machine generated bullet point and categorize each point into one of not helpful, borderline helpful, or very helpful. The highest quality points from the three candidates were combined and refined into a final gold-standard analysis. Concretely, the annotators should first combine all very helpful points, remove duplicate points, reorder the points to maintain coherence, and make necessary textual edits for fluency. Suppose the number of bullet points are lower than our pre-defined lowest threshold (3 bullet points per answer), the annotators should select additional bullet points ranked as borderline helpful to augment the answer. we ask a group of 3 internal members to perform refinement. The agreement accuracy of the refinement process (candidate point selection) is 0.83 and the Cohen’s Kappa is 0.67.
4 Data Statistics
Daco dataset includes training, development and test sets with annotations generated by GPT-4. Furthermore, Daco features a human-refined testing subset. To differentiate the two test sets, we denote the automatically annotated set as TestA and the human-refined set as TestH. Detailed statistics are in Table 1.
| Train | Dev | TestA | TestH | Total | |
| Input Statistics | |||||
| # db | 353 | 22 | 65 | 17 | 440 |
| # queries | 1558 | 100 | 284 | 100 | 1942 |
| Annotation Statistics | |||||
| # bullets | 14.8 | 996 | 2728 | 980 | 19.5 |
| # tokens | 575 | 36.6 | 106 | 42.3 | 760 |
| # code steps | 5086 | 346 | 948 | - | 6380 |
| # code lines | 3.0 | 208 | 555 | - | 3.7 |
| Med. | Max | Min | |
| Detailed statistics for db | |||
| # tables | 1 | 15 | 1 |
| # columns | 15 | 96 | 3 |
| # rows | 572 | 67.2 | 4 |
| Detailed statistics for answer | |||
| # findings | 5 | 8 | 3 |
| # suggestions | 5 | 8 | 3 |
| # tokens | 397 | 864 | 202 |
Database statistics. In total, Daco comprises 440 databases, each of which contains on average 2.3 tables. To better visualize the major topic distribution of this selected subset, we again use BERTopic but group these databases into 10 topics. The keywords for top 5 topics are shown in Figure 4. The leading topic (topic 1) is associated with business setting and consists of 46.52% of the dataset. The remaining nine topics exhibit a relatively even distribution, covering a broad range of domains, including sports (topic 2), healthcare (topic 3), weather (topic 4), and education (topic 5).
| Overlap | % | Representative Examples |
| Low () | 45.6 | Query 1: As an advertising executive, I want to select the channels for targeted ad placements. |
| Query 2: As a platform developer, I want to analyze user behavior and preferences to optimize the user experience. | ||
| Medium () | 52.4 | Query 1: As a farmer, I want to determine the suitable fruit varieties to grow on my farm. |
| Query 2: As a fruit exporter, I want to identify the fruits that meet export standards and have a longer shelf life. | ||
| High () | 2.0 | Query 1: As a consultant for honey market, I want to study the honey production trend to recommend business strategies for my clients. |
| Query 2: As a curious analyst, I want to study the production trend to understand the US honey industry. |
Query statistics. In average, each database is accompanied by 4.4 different user queries. We show the top 15 verbs and their top 3 direct noun objectives in Figure 4. The queries demonstrate a notable level of diversity. The most common type of queries is to request analysis (such as “analyze data” and “identify pattern”), followed by queries aiming to make decisions (such as “determine strategy” and “make decision”). To quantatively verify the diversity of input queries, we measure the overlap between queries over the same database using cosine similarity of Sentence-BERT [32] embeddings. Based on manual inspection of various cases, we categorize the overlap between query pairs into low, medium, and high at thresholds of 0.5 and 0.8. Representative examples are shown in Table 2. We find that 45.6% of generated query pairs have low similarity, 52.4% have medium similarity, and only 2.0% are highly similar. The small percentage of highly similar pairs suggests high diversity of input queries.
Annotation statistics. Figure 7 shows the distribution of the top 10 API functions invoked in the generated code. Excluding the trivial print function, the most frequent APIs are pandas APIs including table manipulation (e.g. groupby and merge) and mathematical computation (e.g. mean and sort_values). Figure 7 and 7 displays the human rating for each bullet points in TestA and TestH collected during the human refinement stage. While automatic annotations significantly lag behind manually refined ones, they still offer valuable knowledge that can be leveraged through supervised fine-tuning, as demonstrated in the experiments section. For manually refined annotations, mentioned in Section 3, the agreement accuracy is 0.83 and Cohen’s Kappa is 0.67, verifying their quality. An example of human refinement is shown in Figure 2.